OpenStudy (anonymous):

If F(x)=f(g(x)), where f(-2)=8, f'(-2)=4, f'(5)=3, g(5)=-2, and g'(5)=6, find F'(5)?

OpenStudy (anonymous):

So if \(F(x)=f(g(x))\) then by chain rule \(F^\prime(x) = f^\prime(g(x))\cdot g^\prime (x)\)

OpenStudy (anonymous):

Or to be specific to this case: \[F^\prime(5) = f^\prime(g(5))\cdot g^\prime (5)\]They had to give you \(f^\prime (-2)\) because \(g(5)=-2\)

OpenStudy (anonymous):

I got it :) its 48

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

I mean 24. lol ;p

OpenStudy (anonymous):

Right